# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ Abstract
In this work, we explore the prospect of generalizing the constant-roll condition in canonical inflationary model to non-canonical models. To find a natural generalization, we focus on three manifestations of this condition and construct constant-roll models corresponding to each manifestation. These models are not equivalent but reduce to the familiar constant-roll model in canonical limit. To showcase the applicability of our generalized mechanism, we examine a specific class of non-canonical models, which can be viewed as extensions of k/G inflation. We conduct a comprehensive analytical examination of the model, elucidating instances where our constant-roll conditions yield disparate outcomes and when they exhibit analogies. We also apply our findings to scrutinize another kinetically driven inflationary model within the constant-roll framework. We demonstrate that each of our constant-roll conditions leads to a unique set of solutions. Afterward, we construct a four-stage constant-roll kinetically driven inflation (or ”constant $k$ inflation” for short) capable of producing primordial black holes, satisfying CMB constraints, and providing a graceful exit from inflation. Employing numerical methods, we analyze this scenario to elucidate how altering the constant-roll condition impacts the power spectrum and the model’s dynamics.
@@ 1 Introduction
It is widely accepted that inflation provides the best explanation for the early universe [1, 2, 3, 4, 5, 6, 7]. It proposes that a small epoch of accelerated expansion at very early universe can potentially smooth out inhomogeneities and anisotropies and reduce the curvature of space. In addition to explaining some of our fundamental questions, such as the flatness, horizon, and monopole problems, it suggests a mechanism for the generation of quantum perturbations in the early universe. There have been countless studies conducted on inflation, and observations of the cosmic microwave background (CMB) and of the large scale structure (LSS) of our universe have confirmed the predictions of inflation. In conventional inflationary models, a scalar field $\phi$ whose potential energy dominates the energy density of the universe is responsible for an accelerated expansion. The spectral index of canonical single field inflationary model is shown to be consistent with observational data in the slow-roll and ultra slow-roll (USR) regimes [8, 9, 10]. Nevertheless, the detailed dynamical characteristics of inflation is still obscure. Motivated by this, a one-parameter class of models (defined by the relation $\ddot{\phi}=\beta H\dot{\phi}$ or equivalently $\ddot{H}=2\beta H\dot{H}$, where $\beta$ is the free parameter of the model) was proposed in [11] to unify simple scenarios like slow-roll and USR, and construct or maybe rule out many models of inflation. Later on the same class of models was studied in more detail and referred to as constant-roll inflation [12]. In the canonical constant-roll scenario, slow-roll behavior can be observed when $\beta\rightarrow 0$ and a constant USR potential at $\beta=-3$. Therefore, it can be considered a generalization of USR inflation.
@@
Although many studies have attempted to generalize the action of the constant-roll model, some authors focused on the generalization of the constant-roll condition [25, 26, 27]. For example, authors in reference [25] add a new free parameter to the constant-roll condition by writing $\ddot{\phi}=\dot{\phi}(\beta H+\alpha)$. Using this condition, we can recover the conventional constant-roll condition for $\alpha=0$, while achieving a broader class of constant-roll potentials for $\alpha\neq 0$.
@@
The generalization of constant-roll condition is much more intriguing in non-canonical models. In a non-canonical model, unlike the canonical case, $\beta=-3$ does not always lead to USR. Therefore, to study the USR limit of the model, an special form of generalized constant-roll condition is required. A good example of such a generalization can be found in [27]. The purpose of their study was to investigate the $f(R)$ gravity under the assumption of a constant rate of roll. For this purpose they defined the constant-roll condition as $\ddot{F}=\beta H_{J}\dot{F}$, where $F=\mathrm{d}f/\mathrm{d}R$ and subscript $J$ denotes the Jordan frame. With this condition, USR and slow-roll regimes amount to $\beta\rightarrow-3$ and $\beta\rightarrow 0$ limits, respectively, and the model is considered to be a natural generalization of USR. The point is that such a generalization can be produced in many ways. One can choose one or another constant-roll condition based on simplicity or aesthetic elegance. In the study of non-canonical models, it is also possible to define an infinite number of distinct constant-roll conditions, all leading to the conventional constant-roll condition $\ddot{\phi}=\beta H\dot{\phi}$ in the canonical limit, yet exhibit varied implications upon departure from the canonical framework. As a result of this freedom, constant-roll condition can be defined in accordance with the needs of the individual to study the desired physical concepts.
@@
In this work, we use different manifestations of constant-roll inflation in order to impose three types of constant-roll conditions on a general $P(X)$ model. Our first approach is based on the fact that the constant-roll condition is defined to ensure a constant rate of field roll. The second approach establishes constant-roll condition that leads to a constant $\epsilon_{2}$, and for our third constant-roll condition, we introduce the condition that corresponds to the USR as $\beta$ approaches $-3$. Next, we present and carefully examine a novel class of constant-roll and propagation inflation (which naturally extends the concept of k/G inflation [67], as it results in k/G inflation when sound speed is one) in order to demonstrate how a model can be affected by changing the constant-roll condition. While each constant-roll condition generally leads to distinct solutions and consequences, we demonstrate that specific limits can be identified where all three conditions converge toward a quasi-canonical phase. We use the term ’quasi-canonical’, in the sense that at these limits slow-roll parameters and consequently the perturbations of the model mimics the canonical model while the action generally differs from the canonical case.
@@
We then broaden our horizons by investigating another model with non-standard kinetic term such as $k$ inflation [68, 69, 70, 71, 72, 73] under the constant-roll assumption. We demonstrate that the model can naturally transition between two phases: the standard canonical constant-roll case and a novel phase we call the large $\gamma$ phase. While the characteristics of the canonical phase remain independent of the specific constant-roll condition employed, the large $\gamma$ phase and the transition dynamics between the two phases are highly dependent on the chosen constant-roll condition. To provide a more vivid demonstration of the constant-roll condition’s impact on the model’s dynamics, we introduce a four-stage constant scenario with the potential to give rise to primordial black holes. We discuss how our scenario aligns with CMB constraints and facilitates a graceful exit from inflation. We conclude that altering the constant-roll condition can modify the resulting power spectrum, thereby significantly influencing primordial black holes abundance.
@@
This paper is arranged as follows. In Section 2, we review the canonical constant-roll model. In Section 3, we introduce three distinct constant-roll conditions to implement constant-roll constraints in non-canonical models. We explore the applications of our generalized constant-roll mechanism by examining a class of non-canonical model in section 4. We carefully scrutinize the interrelationships among our three constant-roll conditions, elucidating their analogies and distinctions. Section 5 investigates another non-canonical model as an example of models that show different dynamical behavior under three defined constant-roll conditions. We then construct a multi-stage constant $k$ scenario capable of producing primordial black holes. Numerical calculations of the power spectrum and other dynamic parameters are presented to understand how different generalizations of the constant-roll condition can impact the model. We conclude in section 6.
@@ 2 Canonical Constant-Roll Inflation
In this section we present a review of the canonical constant-roll model [11, 12, 13, 14, 15]. We determine all possible constant-roll solutions and discuss the primordial perturbations resulting from them. In the next sections, when we discuss more general constant-roll cases, we will follow the same steps as here.
We consider the simplest inflationary scenario that contains a single scalar field with the standard kinetic term minimally coupled to gravity. The dynamics of this model is described by the well-known Friedmann and Klein-Gordon equations [^1] Throughout the paper we work in Planck units where $c=G=\hbar=1$.,
@@
where a comma stands for a partial derivative, and a dot stands for cosmic time derivative. Using these equations, one can find a differential equation in terms of the inflaton field and Hubble parameter
@@
$\epsilon_{1}\equiv\frac{\mathrm{d}\ln H}{\mathrm{~{}d}N},\quad\quad\epsilon_{j+1}\equiv\frac{\mathrm{d}\ln\epsilon_{j}}{\mathrm{~{}d}N},$ (2.4)
where $j$ is a positive integer, and $N$ is the number of e-fold defined as $\mathrm{~{}d}N\equiv\mathrm{d}\ln a=H\mathrm{d}t$. Using definition (2.4) together with equation (2.3) and its derivative, the second slow-roll parameter can be written as
@@
where $\beta$ is a real constant, and for the last two equity we have used equations (2.2) and (2.3). Equation (2.6) is called the constant-roll condition and $\beta$ is the constant-roll parameter. From (2.6) we notice that $\beta$ constrains not only the rate of change of $\dot{H}$ and $\dot{\phi}$, but also the rate of change of potential compared to that of $H^{2}$. Based on equations (2.5) and (2.6), and assuming that $\epsilon_{1}$ is small, $\epsilon_{2}$ can be approximated as $\epsilon_{2}\approx 2\beta$. If we set the parameter $\beta$ to zero, it is similar to the slow-roll case when we neglect the $\ddot{\phi}$ in the Klein-Gordon equation (2.2). If parameter $\beta$ is set to $-3$, it leads to $V_{,\phi}=0$ case which is dubbed as ultra slow-roll inflation (USR) [8, 9, 10]. In canonical models, the three purportedly distinct expressions for $\beta$ are actually ”the same” related by the system of equations (2.1)-(2.3).
In this work, we often find it more convenient to work with number of e-fold $N$ rather than cosmic time $t$. Writing all time derivatives in equations (2.1), (2.2), (2.3), and (2.6) in terms of $N$, we have
@@
By integrating equation (2.10) for the Hubble parameter we find:
@@
where in general $C_{1}$ and $C_{2}$ are two arbitrary complex numbers. However, since the Hubble parameter is real, only real values for $C_{1}$ and $C_{2}$ are physically feasible. For simplicity, we rewrite these constants as $C_{1}=\text{sgn}\left(C_{1}\right)|C_{1}|=\text{sgn}\left(C_{1}\right)M^{2}$ and $C_{2}=\text{sgn}\left(C_{2}\right)|C_{2}|=\text{sgn}\left(C_{2}\right)e^{2\beta N_{0}}$, where $N_{0}$ is a real constant, $M$ is a positive real constant, and sgn is the sign function. From this, we can deduce that the modulus of $C_{2}$ acts as a time shift. Therefore, without loss of generality, we can assume that inflation starts at $N=N_{\text{ini}}$ and set $N_{0}=0$ (or $|C_{2}|=1$). Taking this into account, relation (2.11) can be written as $H^{2}=M^{2}\text{sgn}\left(C_{1}\right)\left(\text{sgn}\left(C_{2}\right)e^{2\beta N}+1\right)$, which leaves us with the following choices:
@@
We first note that the right-hand side of solution (2.12) is always negative, which leads to an imaginary Hubble parameter, and is therefore unacceptable. In contrast, solution (2.13) always results in a real Hubble parameter. For solutions (2.14) and (2.15), the Hubble parameter is real only when we have $\beta N<0$ and $\beta N>0$, respectively. Note that cases (2.14) and (2.15) are zero at $N=0$, which is the singular point of differential equation (2.10). Therefore, our solutions are not viable in the vicinity of this point.
[…]
@@
where $\phi_{c}$ is a constant of integration, and in (2.18) we have used the relation $\mathrm{d}a/\mathrm{d}N=a$. Let us now check whether this solution describes a viable inflationary period. First of all, the parameters $H$, $V$, $\phi$ must be real. For this condition to be satisfied, the constant-roll parameter $\beta$ should be negative (since the field (2.16) is imaginary for $\beta>0$). According to equation (2.8), for a real $\phi_{,N}$, $H_{,N}$ must be negative, which occurs when $\epsilon_{1}$ is positive. Next, we note that inflation is a period of positive acceleration. For acceleration to be positive $\epsilon_{1}$ must be smaller than 1 (this point is easy to understand when one writes $\ddot{a}=aH^{2}(1-\epsilon_{1})$). Lastly, we discuss the potential sign. Using equations (2.1), (2.3), and (2.4), potential can be written as $V/H^{2}=3-\epsilon_{1}$. From this we notice that the potential is positive unless $\epsilon_{1}$ exceeds $3$. But this never happens because inflation already ends when $\epsilon_{1}$ exceeds 1. Putting these together, we conclude that condition $10<\epsilon_{1}<$ 1 is required during inflation. After substituting the Hubble parameter (2.13) into equations (2.4), the slow-roll parameters can be derived as follows:
$\epsilon_{1}(N)=\frac{-\beta}{1+e^{-2\beta N}},\quad\quad\epsilon_{2j}(N)=\frac{2\beta}{1+e^{2\beta N}},\quad\quad\epsilon_{2j+1}(N)=2\epsilon_{1}(N),$ (2.19)
where $j$ is a positive integer. Slow-roll parameters (2.19) are illustrated in the top left panel of Figure 1. In large negative $N$, the first slow-roll parameter can be approximated as $\epsilon_{1}\approx-\beta$, and in large positive $N$ it tends to zero. That means for $-1<\beta<0$ condition $10<\epsilon_{1}<$ 1 is always satisfied. However, for $\beta<-1$, $\epsilon_{1}$ is smaller than 1 only when $2\beta N<\ln\left(-1/(\beta+1)\right)$. Moreover, it is clear that $\beta$ cannot be positive as it leads to negative $\epsilon_{1}$. Consequently, inflation conditions are met when $-1\leq\beta<0$, or when $\beta<-1$ and $2\beta N<\ln\left(-1/(\beta+1)\right)$.
@@
Figure 1: Slow-roll parameters for the canonical constant-roll model. Parameters (2.19), (2.23), (2.27), and (2.34) are plotted in the top left, top right, bottom left, and bottom right panels respectively.
@@
Figure 1 (top right panel) shows these parameters for $\beta=0.5$. In this case, $\epsilon_{1}$ is positive and Hubble parameter is real only if $N<0$ and $\beta>0$. At $N=0$, however, $\epsilon_{1}$ is singular and moves to infinity. Therefore, for the condition $\epsilon_{1}<1$ to be satisfied, $2\beta N$ must remain smaller than $\ln\left(1/(\beta+1)\right)$. For (2.15) case we get
@@
These parameters are plotted in Figure 1 (bottom left panel) for $\beta=-0.5$. Here $\epsilon_{1}$ is positive and Hubble parameter is real only if $N<0$ and $\beta<0$. At large negative $N$ we can approximate $\epsilon_{1}$ to $-\beta$. In this limit condition $10<\epsilon_{1}<$ 1 leads to $-1<\beta<0$. On the other hand, when $N\rightarrow 0$, $\epsilon_{1}$ moves to infinity, and condition $2\beta N>\ln\left(1/(\beta+1)\right)$ is required to keep $\epsilon_{1}$ smaller than $1$.
@@
Heretofore we were able to find all possible constant-roll solutions analytically. However, our solutions are singular at $\beta=0$ (see equations (2.16), (2.20), and (2.24). We can also see the singularity by setting $\beta=0$ in equations (2.13), (2.14), and (2.15). This results in the trivial solution $H=constant$ with $\epsilon_{1}=0$ and a singular $\epsilon_{2}$). To examine the model at this point, we integrate the constant-roll condition (2.10) again, this time by setting $\beta$ to zero. We will find
@@
where, $M^{2}=|C|$ and without losing generality we have set $N_{0}$ to zero. We first note that solution (2.30) is not acceptable as it leads to a positive $H_{,N}$ and an imaginary $\phi_{,N}$. For solution (2.29) the field, potential, and acceleration are as follows:
@@
The slow-roll parameters (2.34) are plotted in the bottom right panel of Figure 1. In this case inflation conditions are satisfied for $N<-1/2$.
With this last solution, we have completely discussed the dynamics of the canonical constant-roll model. We presented our solutions as a functions of $N$ simply because we found it more convenient. Solutions (2.13) and (2.14) correspond to the two cases studied in reference [12], and solution (2.15) corresponds to the special case studied in reference [13]. To our knowledge solution (2.29) has not been discussed previously (probably because it is transient and not dynamically interesting, which we will discuss later).
ptRefer to caption      Refer to captionptRefer to caption    Refer to caption

### v3

@@ Abstract
In this work, we explore the prospect of generalizing the constant-roll condition in canonical inflationary model to non-canonical models. To find a natural generalization, we focus on three manifestations of this condition and construct constant-roll models corresponding to each manifestation. These models are not equivalent but reduce to the familiar constant-roll model in canonical limit. To showcase the applicability of our generalized mechanism, we examine a specific class of non-canonical models, which can be viewed as extensions of k/G inflation. In these models sound speed is constant. We conduct a comparative study, and with an analytical examination of the model, specify instances when our constant-roll conditions yield dissimilar outcomes and when they exhibit analogies. We also apply our findings to scrutinize another kinetically driven inflationary model with varying sound speed. We demonstrate that each of our constant-roll conditions leads to a unique set of solutions. Afterward, we construct a four-stage constant-roll kinetically driven inflation that complies with CMB constraints, it sustains for a sufficiently long period of time, and finally gracefully exits. In this model the spectrum of curvature perturbations is enhanced in a brief phase of non-slow-roll inflationary evolution. Employing numerical methods, we analyse this scenario to elucidate how altering the constant-roll condition impacts the power spectrum and the model’s dynamics.
@@ 1 Introduction
It is widely accepted that inflation provides the best explanation for the early universe [1, 2, 3, 4, 5, 6, 7, 8]. It proposes that a small epoch of accelerated expansion at very early universe can potentially smooth out inhomogeneities and anisotropies and reduce the curvature of space. In addition to explaining some of our fundamental questions, such as the flatness, horizon, and monopole problems, it suggests a mechanism for the generation of quantum perturbations in the early universe. There have been countless studies conducted on inflation, and observations of the cosmic microwave background (CMB) and of the large scale structure (LSS) of our universe have confirmed the predictions of inflation. In conventional inflationary models, a scalar field $\phi$ whose potential energy dominates the energy density of the universe is responsible for an accelerated expansion. The spectral index of canonical single field inflationary model is shown to be consistent with observational data in the slow-roll and ultra slow-roll (USR) regimes [9, 10, 11]. Nevertheless, the detailed dynamical characteristics of inflation is still obscure. Motivated by this, a one-parameter class of models (defined by the relation $\ddot{\phi}=\beta H\dot{\phi}$ or equivalently $\ddot{H}=2\beta H\dot{H}$, where $\beta$ is the free parameter of the model) was proposed in [12] to unify simple scenarios like slow-roll and USR, and construct or maybe rule out many models of inflation. Later on, the same class of models was studied in more detail and referred to as constant-roll inflation in [13]. In the canonical constant-roll scenario, slow-roll behavior can be observed when $\beta\rightarrow 0$ and a constant USR potential at $\beta=-3$. Therefore, it can be considered a generalization of USR inflation.
@@
Although many studies have attempted to generalize the action of the constant-roll model, some authors focused on the generalization of the constant-roll condition [28, 20, 40, 24]. For example, authors in reference [20] add a new free parameter to the constant-roll condition by writing $\ddot{\phi}=\dot{\phi}(\beta H+\alpha)$. Using this condition, the conventional constant-roll is restored with $\alpha=0$, while a broader class of constant-roll potentials is identified for $\alpha\neq 0$.
@@
The generalization of constant-roll condition is much more intriguing in non-canonical models. In a non-canonical model, unlike the canonical case, $\beta=-3$ does not always lead to USR. Therefore, to study the USR limit of the model an special form of generalized constant-roll condition is required. A good example of such a generalization can be found in [28]. The purpose of their study was to investigate the $f(R)$ gravity under the assumption of a constant rate of roll. For this purpose they defined the constant-roll condition as $\ddot{F}=\beta H_{J}\dot{F}$, where $F=\mathrm{d}f/\mathrm{d}R$ and subscript $J$ denotes the Jordan frame. With this condition, USR and slow-roll regimes amount to $\beta\rightarrow-3$ and $\beta\rightarrow 0$ limits, respectively, and the model is considered to be a natural generalization of USR. The point is that such a generalization can be produced in many ways. One can choose one or another constant-roll condition based on simplicity or aesthetic elegance. In the study of non-canonical models, it is also possible to define an infinite number of distinct constant-roll conditions, all leading to the conventional constant-roll condition $\ddot{\phi}=\beta H\dot{\phi}$ in the canonical limit, yet exhibit varied implications upon departure from the canonical framework. As a result of this freedom, constant-roll condition can be defined in accordance with the needs of the individual to study the desired physical concepts.
@@
In this work, we use different manifestations of canonical constant-roll inflation in order to impose three types of constant-roll conditions on a general $P(X,\phi)$ action. Our first approach is based on the fact that the constant-roll condition is defined to ensure a constant rate of roll for the field. The second approach establishes constant-roll condition that leads to a constant $\epsilon_{2}$, and for our third constant-roll condition, we introduce the condition that corresponds to the USR as $\beta$ approaches $-3$. Next, we present and carefully examine a novel class of constant-roll with constant sound speed in order to demonstrate how a model can be affected by changing the constant-roll condition. While each constant-roll condition generally leads to distinct solutions and consequences, we demonstrate that specific limits can be identified where all three conditions converge toward a quasi-canonical phase. We use the term ”quasi-canonical,” in the sense that at these limits slow-roll parameters and consequently the perturbations of the model mimics the canonical model while the action generally differs from the canonical case. A good example in this class of models is k/G inflation [68] in which sound speed is equal to one.
@@
We then broaden our horizons by investigating another model with non-standard kinetic term [69, 70, 71, 72, 73, 74] under the constant-roll assumption. We demonstrate that the model can naturally transition between two phases: a non-canonical constant-roll phase and a canonical phase. While the characteristics of the canonical phase remain independent of the specific constant-roll condition employed, the non-canonical phase and the transition dynamics between the two phases are highly dependent on the chosen constant-roll condition. To provide a more vivid demonstration of the constant-roll condition’s impact on the model’s dynamics, we introduce a four-stage constant $k$ scenario with the potential to give rise to primordial black holes. We discuss how our scenario complies with CMB constraints and facilitates a graceful exit from inflation. We conclude that altering the constant-roll condition can modify the resulting power spectrum, thereby significantly influence on the primordial black holes abundance.
@@
This paper is arranged as follows. In Section 2, we review the canonical constant-roll model. In Section 3, we introduce three distinct constant-roll conditions for non-canonical models. We then explore the applications of our generalized constant-roll mechanism by examining a class of non-canonical model in section 4. We scrutinize the interrelationships in the trio of constant-roll conditions, elucidating their analogies and distinctions. Section 5 investigates another non-canonical model as an example of models that show different dynamical behavior in this comparative study. Next, we construct a multi-stage constant-roll scenario with a brief phase of non-slow-roll inflationary evolution capable of producing primordial black holes. Numerical calculations of the power spectrum and other dynamical parameters are presented to understand how different generalizations of the constant-roll condition can impact the model. We conclude in section 6.
@@ 2 Canonical Constant-Roll Inflation
In this section we present a review of the canonical constant-roll model [12, 13, 15, 17, 21]. We examine the constant-roll solutions and discuss the primordial scalar perturbations generated in these inflationary models. In next sections, when we discuss more general constant-roll cases, we will follow the same steps as here.
We consider the simplest inflationary scenario that contains a single scalar field with the standard kinetic term minimally coupled to gravity. The dynamics of this model is described by the well-known Friedmann and Klein-Gordon equations, [^1] Throughout the paper we work in Planck units where $c=G=\hbar=1$.
@@
where a dot stands for the derivative with respect to cosmic time. Furthermore, we have
@@
$\epsilon_{1}\equiv-\frac{\mathrm{d}\ln H}{\mathrm{~{}d}N},\quad\quad\epsilon_{j+1}\equiv\frac{\mathrm{d}\ln\epsilon_{j}}{\mathrm{~{}d}N},$ (2.4)
where $j$ is a positive integer, and $N$ is the number of $e$-folds defined as $\mathrm{~{}d}N\equiv\mathrm{d}\ln a=H\mathrm{d}t$. Using equation (2.3) and its derivative, it is easy to see that the second slow-roll parameter can be written as
@@
where $\beta$ is a real constant, and for the last two equity we have used equations (2.2) and (2.3). Equation (2.6) is called the constant-roll condition and $\beta$ is the constant-roll parameter. From (2.6) we notice that $\beta$ constrains not only the rate of change of $\dot{H}$ and $\dot{\phi}$, but also the rate of change of potential (compared to that of $H^{2}$). Based on equations (2.5) and (2.6), and assuming that $\epsilon_{1}$ is small, $\epsilon_{2}$ can be approximated as $\epsilon_{2}\approx 2\beta$. This class of models allows to set the parameter $\beta$ to zero to study the standard slow-roll scenario and go beyond to have extremely flat potential $V_{,\phi}=0$ for $\beta=-3$, which is dubbed as ultra slow-roll inflation (USR) [9, 10, 11]. In canonical models, the three purportedly distinct expressions for $\beta$ are actually ’the same’ related by the system of equations (2.1)-(2.3).
In this work, we often find it more convenient to work with number of $e$-folds $N$ rather than cosmic time $t$. Writing all time derivatives in equations (2.1)-(2.3), and (2.6) in terms of $N$, we get
@@
By integrating equation (2.10) for the Hubble parameter we find
@@
where in general $C_{1}$ and $C_{2}$ are two arbitrary complex numbers. However, since the Hubble parameter is real, only real values for $C_{1}$ and $C_{2}$ are physically feasible. For simplicity, we rewrite these constants as $C_{1}=\text{sgn}\left(C_{1}\right)M^{2}$ and $C_{2}=\text{sgn}\left(C_{2}\right)e^{2\beta N_{0}}$, where $N_{0}$ is a real constant, $M$ is a positive real constant, and sgn is the sign function. From this, we can deduce that the modulus of $C_{2}$ acts as a time shift. Therefore, without loss of generality, we can set $N_{0}=0$. Taking this into account, relation (2.11) can be rewritten as $H^{2}=M^{2}\text{sgn}\left(C_{1}\right)\left(\text{sgn}\left(C_{2}\right)e^{2\beta N}+1\right)$, which leaves us with the following choices:
@@
We first note that the right-hand side of solution (2.12) is always negative, which leads to an imaginary Hubble parameter, and is therefore unacceptable. In contrast, solution (2.13) always results in a real Hubble parameter. For solutions (2.14) and (2.15), the Hubble parameter is real only when we have $\beta N<0$ and $\beta N>0$, respectively. Note that cases (2.14) and (2.15) are zero at $N=0$, which is the singular point of differential equation (2.10). Therefore, our solutions are not viable in the vicinity of this point. As we will see later, despite the similarities between justifiable solutions, those describe valid inflationary dynamics for different ranges of $N$ and parameter $\beta$.
@@
where $\phi_{c}$ is a constant of integration, and in (2.18) we have used the relation $\mathrm{d}a/\mathrm{d}N=a$. This case corresponds to the second case studied in [13], as is evident when comparing the potential (2.17) with equation (19) in their work. Let us now check whether this solution describes a viable inflationary period. To have real field, the constant-roll parameter $\beta$ should be negative. According to equation (2.8), for a real $\phi_{,N}$, $H_{,N}$ must be negative, which occurs when $\epsilon_{1}$ is positive, and to have a positive acceleration during inflation, $\epsilon_{1}$ must be smaller than 1 (this point is easy to understand when one writes $\ddot{a}=aH^{2}(1-\epsilon_{1})$). Lastly, using equations (2.1), (2.3), and (2.4), potential can be written as $V/H^{2}=3-\epsilon_{1}$; so the potential is positive unless $\epsilon_{1}$ exceeds $3$. But this never happens with $10<\epsilon_{1}<$ 1.
After substituting the Hubble parameter (2.13) into equations (2.4), the slow-roll parameters can be derived as follows:
$\epsilon_{1}(N)=\frac{-\beta}{1+e^{-2\beta N}},\quad\quad\epsilon_{2j}(N)=\frac{2\beta}{1+e^{2\beta N}},\quad\quad\epsilon_{2j+1}(N)=2\epsilon_{1}(N).$ (2.19)
Slow-roll parameters (2.19) are illustrated in the top left panel of Figure 1 for $\beta=-3$. In large negative $N$, the first slow-roll parameter can be approximated as $\epsilon_{1}\approx-\beta$, and in large positive $N$ it tends to zero. That means for $-1<\beta<0$ condition $10<\epsilon_{1}<$ 1 is always satisfied. However, for $\beta<-1$, $\epsilon_{1}$ is smaller than one only when $2\beta N<\ln\left(-1/(\beta+1)\right)$. Moreover, it is clear that $\beta$ cannot be positive as it leads to negative $\epsilon_{1}$. Consequently, inflation conditions are met for $-1\leq\beta<0$ for all $N$ values, or $\beta<-1$ at $N>\ln\left(-1/(\beta+1)\right)/(2\beta)$.
@@
Figure 1: Slow-roll parameters for the canonical constant-roll model. Parameters (2.19), (2.23), (2.27), and (2.34) are plotted in the top left, top right, bottom left, and bottom right panels, respectively.
@@
Considering potential (2.21), we can readily observe that this case corresponds to the third case studied in the original work [13] (see equation (23) therein). Figure 1 (top right panel) shows parameters (2.23) for $\beta=0.5$. In this case, for the allowed values of $\beta N$, $\epsilon_{1}$ is positive only if $N<0$ and $\beta>0$. At $N=0$, however, $\epsilon_{1}$ is singular and moves to infinity. Therefore, for the condition $\epsilon_{1}<1$ to be satisfied, $2\beta N$ must remain smaller than $\ln\left(1/(\beta+1)\right)$. For (2.15) case we get
@@
Inspection of the potential (2.25) reveals that this final case coincides with the special case analyzed in reference [15]. Parameters (2.27) are plotted in Figure 1 (bottom left panel) for $\beta=-0.5$. Note that for the allowed values of $\beta N$, $\epsilon_{1}$ is positive only if $N<0$ and $\beta<0$. At large negative $N$ we can approximate $\epsilon_{1}$ to $-\beta$. In this limit condition $10<\epsilon_{1}<$ 1 leads to $-1<\beta<0$. On the other hand, when $N\rightarrow 0$, $\epsilon_{1}$ moves to infinity, and condition $2\beta N>\ln\left(1/(\beta+1)\right)$ is required to keep $\epsilon_{1}$ smaller than $1$.
@@
Heretofore, we analysed the general solution of equation (2.10). However, our solutions are singular at $\beta=0$ (see equations (2.16), (2.20), and (2.24)). To examine the model at this parameter value, we integrate the constant-roll condition (2.10) again, this time after setting $\beta$ to zero. We will find
@@
where $M^{2}=|C|$, and without losing generality we have assumed that inflation starts at $N=N_{\text{ini}}$ and have set $N_{0}$ to zero. We first note that solution (2.30) is not acceptable as it leads to a positive $H_{,N}$ and an imaginary $\phi_{,N}$. For solution (2.29) the field, potential, and acceleration are as follows:
@@
The slow-roll parameters (2.34) are plotted in the bottom right panel of Figure 1. In this case inflation conditions are satisfied for $N<-1/2$. With this last solution, we have completely discussed the dynamics of the canonical constant-roll model.
Refer to caption      Refer to captionRefer to caption    Refer to caption